# Oracle-based Uniform Sampling from Convex Bodies

### v1

@@ 2.2 Alternating sampling framework
While [42] proposes ASF for log-concave sampling, [8] extends the assumption of log-concave distributions to distributions satisfying common isoperimetric inequalities such as Log-Sobolev inequality or Poincaré inequality. The crucial observation by [8] is that while ASF is a Markov chain, each iteration of this chain can be viewed as a pair of forward and backward diffusion steps where probabilistic tools for Itô diffusion processes can be applied. For the ASF, i.e., Algorithm 1, the first step is generating a Gaussian sample and thus can be easily done, while the second step is non-trivial and is called the Restricted Gaussian Oracle (RGO). In both [42] and [8], the authors either assume they have exact access to the RGO, or that $f$ is smooth so that RGO can be easily done via rejection sampling. Novel realizations of the RGO to either reduce its cost or to relax the smoothness assumption have been investigated in [48, 22, 18, 65, 49, 50] among others.
The uniform distribution on $K$ has density proportional to $\mathbf{1}_{K}$, where $\mathbf{1}_{K}(x)$ equals $1$ on $K$ and 0 otherwise. In the context of the ASF introduced above, if we take $f(x)=I_{K}(x)$ where the indicator function $I_{K}(x)$ equals 0 if $x\in K$ and equals $\infty$ otherwise, then the target of the ASF will be the uniform distribution on $K$. In particular, we have
@@
Moreover, denote ${\cal N}(y,\eta I_{d})|_{K}$ the Gaussian distribution ${\cal N}(y_{k},\eta I_{d})$ restricted to $K$, i.e.,
@@
Then, Algorithm 1 for uniform sampling on $K$ turns into
Algorithm 2 ASF for the uniform distribution on $K$
@@
The upcoming result is about contractivity in Rényi divergence and $\chi^{2}$-divergence of Algorithm 2. Via a clever smoothing argument, the authors of [39] are able to adapt the proof technique in [8] to the case of uniform sampling from convex $K$. In particular, as pointed out in [39], the upcoming result does not require convexity of $K$.
@@
([39, Theorem 23]) Let $\mu^{X}_{k}$ be the law of the $k$-th output of Algorithm 2 (ASF). Denote $C_{\mathrm{PI}}$ and $C_{\mathrm{LSI}}$ respectively the Poincaré constant and the log Sobolev constant of the uniform distribution $\pi^{X}$ on $K$ whose asymptotics are provided in Lemma 2.1. Then for any $q\geq 1$,
Per Theorem 2.2, we can immediately deduce the iteration complexity of Algorithm 2 under the assumption that the RGO implementation has no cost. The proof is deferred to Appendix A. Note that we assume $K$ is convex in the upcoming result to be able to cite known asymptotics of the PI constant and the LSI constant of the uniform distribution on $K$ (Lemma 2.1).
Corollary 2.3
Let $K\subset\mathbb{R}^{d}$ be a convex set. Assume Algorithm 2 starts from an $M$-warm distribution $\mu^{X}_{0}$, i.e., $\frac{d\mu_{0}}{d\pi^{X}}\leq M$. Let $\epsilon>0$. Denote $C_{\mathrm{PI}}$ and $C_{\mathrm{LSI}}$ respectively the Poincaré constant and the log Sobolev constant of the uniform distribution $\pi^{X}$ on $K$. Then,
@@
    with respect to the Rényi divergence $\mathcal{R}_{q}$ and $q\geq 1$, the algorithm can achieve $\epsilon$-accuracy within
    $k^{*}=O\left(d^{2}{C_{\mathrm{LSI}}q\log\left(2\frac{\log M}{\epsilon}\right)}\right)$ (5)
    iterations, where, in view of Lemma 2.1, $C_{\mathrm{LSI}}=\mathcal{O}(D^{2})$ in general and $C_{\mathrm{LSI}}=\mathcal{O}(D)$ if $\pi^{X}$ is isotropic;
@@
    with respect to the $\chi^{2}$-divergence, the algorithm can achieve $\epsilon$-accuracy within
    $k^{**}=O\left(d^{2}{C_{\mathrm{PI}}\log\left(2\frac{M^{2}+1}{\epsilon}\right)}\right)$ (6)
    iterations, where, in view of Lemma 2.1, $C_{\mathrm{PI}}(\pi)=\mathcal{O}\left(\left\lVert\mathrm{Cov}{\left(\pi\right)}\right\rVert_{\operatorname{op}}\log d\right)$ in general and $C_{\mathrm{PI}}(\pi)=\mathcal{O}\left(\log d\right)$ if $\pi^{X}$ is isotropic.
The iteration complexities provided in Corollary 2.3 have not taken into account methods to implement the RGO (Step 2 in Algorithm 2) and the costs associated with them. Our method for the RGO implementation in the upcoming sections is based on rejection sampling, which requires the construction of a suitable sampling proposal that is close to the target $\pi^{X|Y}(x|y)={\cal N}(y,\eta I_{d})|_{K}$ in Algorithm 2. We will see that one can construct quite natural proposals if given access to either a projection oracle or a separation oracle on $K$.
Remark 2.4
We observe here a few basic facts. First, via (2) and Lemma A.1(a), one immediately gets
[…]
Furthermore, the warm start condition (A2) implies that the same warmness holds between $\pi^{Y}$ and $\mu_{\eta}$, which is the output of $\mu_{0}$ after the first step of Algorithm 2, i.e.,
$\frac{\mathrm{d}\mu_{\eta}}{\mathrm{d}\pi^{Y}}\leq M.$ (8)
Indeed, assume any $U\subseteq\mathbb{R}^{d}$. For $y\in\mathbb{R}^{d}$, set $U-y:=\{x\in\mathbb{R}^{d}:x+y\in U\}$. Denote $\gamma(\cdot)$ the density of $\mathcal{N}(0,\eta I_{d})$. Then, by $\mathrm{d}\mu_{0}/\mathrm{d}\pi^{X}\leq M$, $\mu_{\eta}=\mu_{0}*\gamma$, and $\pi^{Y}=\pi^{X}*\gamma$, we have
$\mu_{\eta}(U)=\int_{\mathbb{R}^{d}}\mu_{0}(U-y)\gamma(y)\mathrm{d}y$ $=\int_{\mathbb{R}^{d}}\int_{U-y}\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\pi^{X}}(t)\mathrm{d}\pi^{X}(t)\gamma(y)\mathrm{d}y$
$\leq M\int_{\mathbb{R}^{d}}\pi^{X}(U-y)\gamma(y)\mathrm{d}y=M\pi^{Y}(U).$
Consequently, this shows (8) holds, and further the warmness holds for every step of Algorithm 2.
@@ 3 Projection oracle-based proximal sampling
This section aims to implement the RGO step, i.e., Step 2 in Algorithm 2, via rejection sampling and the projection oracle of $K$. At the $k$-th iteration, Step 1 in Algorithm 2 generates from ${\cal N}(x_{k},\eta I_{d})$ a point $y:=y_{k}$, which is fixed in RGO. Then the RGO step is supposed to sample from the truncated Gaussian ${\cal N}(y,\eta I_{d})|_{K}$. To implement RGO by rejection sampling, we need to construct a proposal that is both easier to sample than ${\cal N}(y,\eta I_{d})|_{K}$ and also reasonably close to ${\cal N}(y,\eta I_{d})|_{K}$ in order to ensure the acceptance probability is high, or equivalently, the number of rejections is low. Examining the equivalent formulas of ${\cal N}(y,\eta I_{d})|_{K}$ in (3), one can easily figure out that $\mathcal{N}(y,\eta I_{d})|_{K}$ concentrates at
@@
Below is our implementation of RGO via the projection oracle $\operatorname{proj}_{K}$ and rejection sampling. ${\cal U}[0,1]$ will denote the uniform distribution on $[0,1]$.
@@
Per Lemma A.3 (which can be found in Appendix A), the sample $X$ generated by Algorithm 3 follows the distribution $\pi^{X|Y}$, and therefore our RGO implementation is unbiased. Moreover, the acceptance condition (10) in Algorithm 3 guarantees that the output of the algorithm is in $K$ almost surely (i.e., the output is a feasible point). Indeed, when $X\notin K$, $\mathbf{1}_{K}(X)=0$ and the right hand side of (10) equals 0, so that the probability that $X$ is not in $K$ while being accepted equals to $\Pr\left(U\leq 0\right)=0$. We will see later in Section 4 another RGO implementation via a separation oracle on $K$ and rejection sampling that also guarantees the accepted $X\in K$ almost surely.
As a comparison, the In-and-Out algorithm by [39] samples $x\sim\pi^{X|Y}(x|y)={\cal N}(y,\eta I_{d})|_{K}$ by sampling $x_{i}\sim\mathcal{N}(y,\eta I_{d})$ up to $N=\tilde{\cal O}(d^{2})$ times (see [39, Remark 2]) until one finds a point $x_{i}\in K$; otherwise the algorithm stops and declares failure. While this only requires a membership oracle, repeatedly sampling $x_{i}\sim\mathcal{N}(y,\eta I_{d})$ does not encourage the desired scenario that $x_{i}\in K$ for some $i\leq N$, and thus some failure probability of their algorithm is to be expected. We note that the aforementioned failure probability is removed in [34] with an introduction of the restart procedure therein. Here, we are offering an alternative to the restart procedure by using either a projection oracle or a separation oracle on $K$.
Next, we want to make sure the right hand side of (10) is no more than $1$ for the acceptance/rejection condition (10) to be well-defined. In addition, we also introduce a new function $\mathcal{P}_{1}$ that will naturally appear later in the analysis of the rejection sampling.
@@
For every $x\in\mathbb{R}^{d}$, we have
$I_{K}(x)-\frac{1}{\eta}\langle x-\operatorname{proj}_{K}(y)\,,\operatorname{proj}_{K}(y)-y\rangle\leq 0,$ (11)
and hence the acceptance test (10) is well-defined. Moreover, (10) is equivalent to
$U\leq\exp(\mathcal{P}_{1}(x)-\Theta_{y}^{\eta,K}(x)),$ (12)
where $\Theta_{y}^{\eta,K}$ is as in (9) and
@@
Proof: It follows from the convexity of $K$ that $\langle x-\operatorname{proj}_{K}(y)\,,\operatorname{proj}_{K}(y)-y\rangle\geq 0$, and hence that (11) holds. In view of (9) and (13), we observe that the RHS of (10) is equivalent to
Hence, the proof is completed.
The following lemma is one of the key technical contributions of the paper. It will allow us to bound the average number of rejections of Algorithm 3 in Proposition 3.4, and also the average number of rejections of Algorithm 4 in Proposition 4.2.
@@
It follows from $B(0,1)\subseteq K$ in condition (A1) that $K_{\delta}=K+\delta B(0,1)\subseteq(1+\delta)K$. This relation, the fact that $(1+\delta)^{d-1}\leq\exp(\delta d)$, and Lemma A.2 together imply that
@@
Plugging this inequality into (15), we obtain
$\int_{K^{c}}\exp\left(-\frac{\left(\|\operatorname{proj}_{K}(y)-y\|-\tau\right)^{2}}{2\eta}\right)\mathrm{d}y\leq d\mathsf{vol}(K)\int_{0}^{\infty}\exp\left(-\frac{\left(\delta-\tau\right)^{2}}{2\eta}+\delta d\right)\mathrm{d}\delta$
@@
$=$ $d\mathsf{vol}(K)\exp\left(\frac{\eta d^{2}}{2}+\tau d\right)\int_{-\eta d-\tau}^{\infty}\exp\left(-\frac{a^{2}}{2\eta}\right)\mathrm{d}a\leq\mathsf{vol}(K)\exp\left(\frac{\eta d^{2}}{2}+\tau d\right)\sqrt{2\pi\eta d^{2}},$
where $a=\delta-\tau-\eta d$. The lemma finally follows from combining the above inequality and (14).
Next, we are able to deduce the average number of rejections in Algorithm 3 for each iteration of the ASF (Algorithm 2). Our step size $\eta=1/d^{2}$ implies that the average numbers of rejections is of the order $\mathcal{O}(1)$ and matches the dimension dependence of the chosen step size in [39, Theorem 27].
Proposition 3.4
Assume conditions (A1) and (A2) hold, and consider Algorithm 2 with step size $\eta=1/d^{2}$. Then, the average number of rejections in Algorithm 3 is bounded by $M(\sqrt{2\pi e}+1)$.
Proof: Denote $\mu_{\eta}$ the distribution of $y=y_{k}$ for the first step of Algorithm 2. Per Lemma A.3, the average number of rejections is $\mathbb{E}_{\mu_{\eta}}[n_{y}]$ where $n_{y}$ is defined in (21). The fact that $d\mu_{\eta}/d\pi^{Y}\leq M$ from (8) implies
$\mathbb{E}_{\mu_{\eta}}[n_{y}]\leq M\mathbb{E}_{\pi^{Y}}[n_{y}],$
$\mathbb{E}_{\pi^{Y}}[n_{y}]=\int_{\mathbb{R}^{d}}\frac{\int_{\mathbb{R}^{d}}\exp(-\mathcal{P}_{1}(x))\mathrm{d}x}{\int_{K}\exp\left(-\frac{1}{2\eta}\|x-y\|^{2}\right)\mathrm{d}x}\pi^{Y}(y)\mathrm{d}y.$
Via Lemma A.1(a), it is easy to compute that
@@
The above two identities and (7) yield
$\mathbb{E}_{\pi^{Y}}[n_{y}]=\frac{\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2\eta}\|\operatorname{proj}_{K}(y)-y\|^{2}\right)\mathrm{d}y}{\mathsf{vol}(K)}.$
Finally, it follows from Lemma 3.3 with $\tau=0$ that
$\mathbb{E}_{\pi^{Y}}[n_{y}]\leq\exp\left(\frac{\eta d^{2}}{2}\right)\sqrt{2\pi\eta d^{2}}+1\quad\text{and}\quad\mathbb{E}_{\mu_{\eta}}[n_{y}]\leq M\exp\left(\frac{\eta d^{2}}{2}\right)\sqrt{2\pi\eta d^{2}}+M.$
@@
Finally, we are ready to present the main result of Section 3 by putting together Corollary 2.3 and Proposition  3.4.

### v2

@@ 2.2 The proximal sampler
While [48] proposes the proximal sampler for log-concave sampling, [11] extends the assumption of log-concave distributions to distributions satisfying common isoperimetric inequalities such as log-Sobolev inequality or Poincaré inequality. The crucial observation by [11] is that while the proximal sampler is a Markov chain, each iteration of this chain can be viewed as a pair of forward and backward diffusion steps where probabilistic tools for Itô diffusion processes can be applied. For the proximal sampler, i.e., Algorithm 1, the first step is generating a Gaussian sample and thus can be easily done, while the second step is non-trivial RGO. In both [48, 11], the authors either assume they have exact access to the RGO, or that $f$ is smooth so that RGO can be easily realized via rejection sampling. Novel realizations of the RGO to either reduce its cost or to relax the smoothness assumption have been investigated in [55, 29, 24, 76, 56, 57] among others.
The uniform distribution on $K$ has density proportional to $\mathbf{1}_{K}$, where $\mathbf{1}_{K}(x)$ equals $1$ on $K$ and 0 otherwise. In the context of the proximal sampler introduced above, if we take $f(x)=I_{K}(x)$ where the indicator function $I_{K}(x)$ equals 0 if $x\in K$ and equals $\infty$ otherwise, then the target of the proximal sampler will be the uniform distribution on $K$. In particular, we have
@@
Moreover, denote ${\cal N}(y,\eta I_{d})|_{K}$ the Gaussian distribution ${\cal N}(y,\eta I_{d})$ restricted to $K$, i.e.,
@@
Then, Algorithm 1 for uniform sampling on $K$ turns into the following proximal sampler.
Algorithm 2 Proximal sampler for the uniform distribution on $K$
@@
The next theorem follows directly from the one-step contraction in $\chi^{2}$ and Rényi divergences proved in [44, Theorem 23] (restated as Theorem B.6 in Appendix B). Assuming the RGO step (Algorithm 2, Step 2) is available and exact, we obtain the iteration complexity of the proximal sampler in the next theorem. The proof is in Appendix B.
@@
Let $K\subset\mathbb{R}^{d}$ be a convex set. Assume Algorithm 2 starts from an $M$-warm distribution $\mu^{X}_{0}$, i.e., $\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\pi^{X}}\leq M$. Let $\epsilon>0$. Denote $C_{\mathrm{PI}}$ and $C_{\mathrm{LSI}}$ respectively the Poincaré constant and the log Sobolev constant of the uniform distribution $\pi^{X}$ on $K$. Then,
@@
    with respect to the Rényi divergence $\mathcal{R}_{q}$ and $q\geq 1$, the algorithm can achieve $\epsilon$-accuracy within $O\left(\frac{1}{\eta}{C_{\mathrm{LSI}}q\log\left(2\frac{\log M}{\epsilon}\right)}\right)$ iterations, where, per Lemma 2.1, $C_{\mathrm{LSI}}=\mathcal{O}(D^{2})$ in general and $C_{\mathrm{LSI}}=\mathcal{O}(D)$ if $\pi^{X}$ is isotropic;
@@
    with respect to the $\chi^{2}$-divergence, the algorithm reaches $\epsilon$-accuracy in $O\left(\frac{1}{\eta}{C_{\mathrm{PI}}\log\left(2\frac{M^{2}+1}{\epsilon}\right)}\right)$ iterations, where, per Lemma 2.1, $C_{\mathrm{PI}}(\pi)=\mathcal{O}\left(\left\lVert\mathrm{Cov}{\left(\pi\right)}\right\rVert_{\operatorname{op}}\log d\right)$ in general and $C_{\mathrm{PI}}(\pi)=\mathcal{O}\left(\log d\right)$ if $\pi^{X}$ is isotropic.
The iteration complexities provided in Theorem 2.2 have not taken into account methods to implement the RGO (Step 2 in Algorithm 2) and the costs associated with them. Our method for the RGO implementation in the upcoming sections is based on rejection sampling, which requires the construction of a suitable sampling proposal that is close to the target $\pi^{X|Y}(x|y)={\cal N}(y,\eta I_{d})|_{K}$ in Algorithm 2. We will see that one can construct quite natural proposals if given access to either a projection oracle or a separation oracle on $K$.
@@ 3 Projection oracle-based proximal sampling
This section aims to implement the RGO step, i.e., Step 2 in Algorithm 2, via rejection sampling and the projection oracle of $K$. At the $k$-th iteration, Step 1 in Algorithm 2 generates from ${\cal N}(x_{k},\eta I_{d})$ a point $y:=y_{k}$, which is fixed in RGO. Then the RGO step is supposed to sample from the truncated Gaussian ${\cal N}(y,\eta I_{d})|_{K}$. To implement RGO by rejection sampling, we need to construct a proposal that is both easier to sample than ${\cal N}(y,\eta I_{d})|_{K}$ and also reasonably close to ${\cal N}(y,\eta I_{d})|_{K}$ in order to ensure the acceptance probability is high, or equivalently, the number of proposals is low. Examining the equivalent formulas of ${\cal N}(y,\eta I_{d})|_{K}$ in (3), one can easily figure out that $\mathcal{N}(y,\eta I_{d})|_{K}$ concentrates at
@@
Below is our implementation of RGO via the projection oracle $\operatorname{proj}_{K}$ and rejection sampling. Let ${\cal U}[0,1]$ denote the uniform distribution on $[0,1]$.
@@
By Lemma B.3 (Appendix B), the output $X$ of Algorithm 3 has law $\pi^{X|Y}$, so the RGO implementation is unbiased. Moreover, the acceptance test (5) enforces feasibility: if $X\notin K$ then $\mathbf{1}_{K}(X)=0$, hence $\Pr(\text{$X$ accepted and }X\notin K)=\Pr(U\leq 0)=0$. Section 4 gives an alternative RGO implementation via a separation oracle that also accepts only feasible points.
In contrast, [41] draws $x_{i}\sim{\cal N}(y,\eta I_{d})$ up to $\tilde{\cal O}(d^{2})$ times and accepts the first $x_{i}\in K$, declaring failure otherwise [41, Remark 2]. This membership-oracle approach has nonzero failure probability, noting that this failure is removed by [45] via a restart procedure. Here, we are offering an alternative to the restart procedure by using a projection oracle or a separation oracle on $K$.
In the next lemma (proved in Appendix B), we verify that the acceptance test (5) is well-defined. We also introduce a function $\mathcal{P}_{1}$, which will appear naturally in later rejection-sampling analysis.
@@
Assume condition (A1) holds. Then for every $x\in\mathbb{R}^{d}$, we have
$I_{K}(x)-\frac{1}{\eta}\langle x-\operatorname{proj}_{K}(y)\,,\operatorname{proj}_{K}(y)-y\rangle\leq 0$
and hence the acceptance test (5) is well-defined. Moreover, (5) is equivalent to $U\leq\exp(\mathcal{P}_{1}(x)-\Theta_{y}^{\eta,K}(x))$ where $\Theta_{y}^{\eta,K}$ is as in (4) and
@@
The following lemma is a key technical contribution. It enables the bounds on the average number of proposals of Algorithm 3 in Theorem 3.5 and of Algorithm 4 in Theorem 4.2.
@@
It follows from $B(0,1)\subseteq K$ in condition (A1) that $K_{\delta}=K+\delta B(0,1)\subseteq(1+\delta)K$. This relation, the fact that $(1+\delta)^{d}\leq\exp(\delta d)$, and Lemma B.2 with $K=K_{\delta}$ together imply that
@@
Plugging this inequality (9) into (8), we obtain
[…]
@@
$=$ $d\mathsf{vol}(K)\exp\left(\frac{\eta d^{2}}{2}+\tau d\right)\int_{-\eta d-\tau}^{\infty}\exp\left(-\frac{a^{2}}{2\eta}\right)\mathrm{d}a\leq\mathsf{vol}(K)\exp\left(\frac{\eta d^{2}}{2}+\tau d\right)\sqrt{2\pi\eta d^{2}},$ (10)
where $a=\delta-\tau-\eta d$. The stated lemma finally follows from combining (3) and (7).
Remark 3.4
Regarding the proof argument above, we provide in Lemma B.5 an alternative way of bounding the terms in (8), which uses integration by parts and the fact that $\frac{\mathrm{d}}{\mathrm{d}\delta}\mathsf{vol}(K_{\delta})=\mathsf{vol}_{d-1}(\partial K_{\delta})$ for $\delta\geq 0$ a.s..
Next, we bound the average number of proposals in Algorithm 3 per each iteration of Algorithm 2. With the stepsize $\eta=1/d^{2}$, this average is $\mathcal{O}(1)$ and matches the dimension scaling in [41, Theorem 27].
Theorem 3.5
Assume conditions (A1) and (A2) hold, and consider Algorithm 2 with stepsize $\eta=1/d^{2}$. Then, the average number of proposals in Algorithm 3 is bounded by $M(\sqrt{2\pi e}+1)$.
Proof: Denote $\mu^{k}$ the distribution of $y=y_{k}$ for the first step of Algorithm 2. Write $n_{y}$ the average number of proposals conditioned on $y$ in the rejection sampler that is Algorithm 3. Per Lemma B.3, the average number of proposals is $\mathbb{E}_{\mu^{k}}[n_{y}]$ where the formula of $n_{y}$ is given in (16). The fact that $d\mu^{k}/d\pi^{Y}\leq M$ from Lemma B.4 implies $\mathbb{E}_{\mu^{k}}[n_{y}]\leq M\mathbb{E}_{\pi^{Y}}[n_{y}]$, and hence we will focus on bounding $\mathbb{E}_{\pi^{Y}}[n_{y}]$. In view of (16), the latter expression becomes
$\mathbb{E}_{\pi^{Y}}[n_{y}]=\mathbb{E}_{\pi^{Y}}\left[\frac{N_{\mathcal{P}_{1}}}{N_{\Theta_{y}^{\eta,K}}}\right]=\int_{\mathbb{R}^{d}}\frac{\int_{\mathbb{R}^{d}}\exp(-\mathcal{P}_{1}(x))\mathrm{d}x}{\int_{K}\exp\left(-\frac{1}{2\eta}\|x-y\|^{2}\right)\mathrm{d}x}\pi^{Y}(y)\mathrm{d}y.$
Via Lemma B.1(a), it is easy to get
@@
Furthermore, via (2) and Lemma B.1(a), one obtains
[…]
The last three identities yield $\mathbb{E}_{\pi^{Y}}[n_{y}]=\frac{1}{\mathsf{vol}(K)}\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2\eta}\|\operatorname{proj}_{K}(y)-y\|^{2}\right)\mathrm{d}y$. Finally, it follows from Lemma 3.3 with $\tau=0$ that
$\mathbb{E}_{\pi^{Y}}[n_{y}]\leq\exp\left(\frac{\eta d^{2}}{2}\right)\sqrt{2\pi\eta d^{2}}+1\quad\text{and}\quad\mathbb{E}_{\mu^{k}}[n_{y}]\leq M\exp\left(\frac{\eta d^{2}}{2}\right)\sqrt{2\pi\eta d^{2}}+M.$
@@
4 Separation oracle-based proximal sampling
In Section 3, when a projection oracle is available, one get an exact solution to $\mathrm{argmin}\,_{x\in\mathbb{R}^{d}}\Theta^{\eta,K}_{y}(x)$ as $\operatorname{proj}_{K}(y)$, and the Gaussian proposal for the rejection sampling can thus be centered at $\operatorname{proj}_{K}(y)$. When $\operatorname{proj}_{K}$ is not available, we propose to use a state-of-the-art Cutting Plane method by [33], which uses a separation oracle on $K$ to find an approximate solution $\hat{x}$ of $\mathrm{argmin}\,_{x\in\mathbb{R}^{d}}\Theta^{\eta,K}_{y}(x)$. At this point, using $\hat{x}$, we do rejection sampling with the proposal
[…]
As in Section 3, this proposal is centered at a high-concentration point $\hat{x}$, but it is not Gaussian. Sampling from $\nu$ is nevertheless simple: it reduces to a one-dimensional sampling problem (see Lemma C.3 in Appendix C).
Below are our RGO implementation assuming access to a separation oracle on $K$.
Algorithm 4 Separation oracle-based implementation of RGO
1. Compute a $(1/d)$-solution $\hat{x}$ of $\mathrm{argmin}\,_{x\in\mathbb{R}^{d}}\Theta^{\eta,K}_{y}(x)$ via the Cutting Plane method by [33].
2. Generate $U\sim{\cal U}[0,1]$ and $X\sim\nu(x)$ via Algorithm 5 in Appendix C.4.
3. If
$U\leq\exp\left(\mathcal{P}_{2}(X)-\Theta_{y}^{\eta,K}(X)\right),$ (12)
then accept $X$; otherwise, reject $X$ and go to step 2. The function $\mathcal{P}_{2}$ is defined in (13) below.
Lemma C.2 in Appendix C.1 bounds the number of separation-oracle calls needed to produce $\hat{x}$ in Step 1. Lemma B.3 in Appendix B shows that the rejection sampling is unbiased.
The following result (proved in Appendix C.2) ensure the acceptance test at (12) is well-defined.
Lemma 4.1
Assume condition (A1) holds. Recall $\Theta^{\eta,K}_{y}$ and $\mathcal{P}_{1}$ defined in (4) and (6), respectively. Set
Then, for every $x\in\mathbb{R}^{d}$, $\Theta^{\eta,K}_{y}(x)\geq\mathcal{P}_{1}(x)\geq\mathcal{P}_{2}(x)$. In particular, the fact that $\Theta^{\eta,K}_{y}\geq\mathcal{P}_{2}$ ensures the acceptance test (12) is well-defined.
Next, we bound the average number of proposals and the number of separation-oracle queries in Algorithm 4 per each iteration of Algorithm 2. The proof is deferred to Appendix C.3.